# Query Routing for Retrieval-Augmented Language Models

### v1

@@ Appendix C Routing Performance under Noisy Retrieval
Appendix D Full Results on Latency-Aware Routing
PopQA (Local) PopQA (Online) NQ WQ
RouterDC 44.58 47.04 0.02 49.80 51.85 - 33.10 42.92 - 46.39 49.17 -
MF - 46.30 - 49.96 52.22 -0.91 37.34 44.17 - - 50.42 1.66
GraphRouter - 47.41 -0.50 - 51.48 - 35.68 38.35 0.02 45.67 48.18 0.01
KNN Router - 46.67 -0.06 - 52.22 - 41.13 50.63 0.72 - 50.42 1.72
Appendix E Full Results of Sensitivity Analysis and Ablation Study
Figure 9: Effects of dimension (baseline: 128).
Figure 10: Effects of $\lambda$ (baseline: $\lambda=0$).
Table 9 presents the impact of the dimensionality of LLM knowledge embeddings and RAG capability embeddings on test accuracy in the RAGRouter architecture, with Figure 10 showing testing accuracy change rates relative to the 128-dimensional baseline. The best average accuracy is observed at dimension 768.
Table 9: Effects of dimension.
PopQA MedMCQA
Dimension Local Online Local Online NQ WQ TQA Avg
128 48.52 52.22 71.11 73.70 54.17 53.33 87.92 63.00
256 48.52 53.33 70.37 73.33 55.42 55.83 88.75 63.65
384 49.26 52.59 70.37 72.59 56.25 53.75 89.17 63.43
512 48.89 52.59 70.37 73.33 55.42 56.25 89.17 63.72
768 48.52 52.59 71.48 74.44 56.67 56.67 90.83 64.46
1024 48.52 52.59 71.11 73.70 55.42 51.67 90.00 63.29
Table 10 presents the impact of the loss weight $\lambda$ on test accuracy, with Figure 10 showing testing accuracy change rates relative to the baseline of $\lambda=0$. The best average accuracy is observed at $\lambda=2$.
Table 10: Effects of $\lambda$.
PopQA MedMCQA
$\lambda$ Local Online Local Online NQ WQ TQA Avg
0.0 48.89 52.59 70.37 74.44 55.42 54.17 90.42 63.76
0.5 48.89 52.96 70.74 74.07 55.42 54.17 90.42 63.81
1.0 48.52 52.59 71.48 73.70 55.42 55.42 90.42 63.94
1.5 48.89 52.59 71.48 74.07 55.83 55.42 90.83 64.16
2.0 48.52 52.59 71.48 74.44 56.67 56.67 90.83 64.46
2.5 48.52 52.96 71.48 74.44 57.08 56.25 90.42 64.45
3.0 48.52 52.96 71.85 74.44 55.83 55.83 90.00 64.21
Appendix F Case Study
Query What is Agen the capital of?
Ground truth Lot-et-Garonne
Router RAGRouter MF
Selected LLM Qwen2.5-14B-Instruct Meta-Llama-3.1-8B-Instruct
Response w/o RAG Lot department. () Lot-et-Garonne department. (✓)
Response w/ RAG Lot-et-Garonne department. (✓) The prune capital. ()
Query who are the dancers in the lazy song video?
Ground truth Poreotics
Router RAGRouter RouterDC
Selected LLM Llama-3.3-70B-Instruct Qwen2.5-72B-Instruct
Response w/o RAG Bruno Mars and dancers. () Five dancers. ()
Appendix G Inference Cost of RAGRouter
We report the inference characteristics of RAGRouter on a single NVIDIA RTX 4090D GPU. The peak GPU memory usage is 4147 MiB. With a batch size of 64, RAGRouter processes 270 instances in 3 seconds over 5 batches, resulting in an average inference time of 0.011 seconds per instance. These results demonstrate that RAGRouter is both lightweight and efficient during inference.
Appendix H Limitations and Future Works
Beyond RAG as the mainstream paradigm, recent techniques such as fine-tuning [20, 19], knowledge distillation [18], and Chain-of-Thought prompting [51] have also been widely adopted to enhance the performance of LLM. These approaches inherently involve explicit or implicit modulation of model knowledge, often entailing complex processes of knowledge transfer or shift. Although this work focuses on capability modeling and dynamic routing within RAG settings, we believe the proposed framework and analytical perspective offer valuable insights into the representation and utilization of LLM capabilities in broader contexts. As this lies beyond the scope of this paper, we leave the exploration to future work.

### v2

@@ Appendix G Routing Performance under Noisy Retrieval
Appendix H Routing Performance on Summarization Task
To verify that RAGRouter is also applicable to RAG tasks beyond question answering, we conduct experiments on WikiASP [16], a summarization task commonly used in RAG-related research. Specifically, WikiASP aims to generate aspect-based summaries for Wikipedia entities and naturally supports the use of retrieved evidence. We adopt ROUGE-L as the evaluation metric.
For continuous performance metrics such as ROUGE-L, we extend RAGRouter’s contrastive learning by score-based probabilistic sampling. Given each candidate LLM’s normalized score $s\in[0,1]$, we label it as “can answer” (positive, label 1) with probability $s$, and “cannot answer” (negative, label 0) with probability 1s1-s. This strategy is applied in our WikiASP experiments, enabling contrastive training based on continuous performance signals rather than discrete correctness labels.
Table 17: Performance of RAGRouter and baselines on WikiASP measured by ROUGE-L.
Method WikiASP
Qwen2.5-0.5B-Instruct 0.1388
Qwen2.5-1.5B-Instruct 0.1411
Llama-3.2-3B-Instruct 0.1768
Qwen2.5-3B-Instruct 0.1528
Qwen2.5-7B-Instruct 0.1897
Llama-3.1-8B-Instruct 0.1455
KNN Router 0.1881
MF 0.1865
RAGRouter 0.1981
As shown in Table 17, RAGRouter outperforms non-RAG-aware routing baselines by more than 5.3%. These findings demonstrate that RAGRouter generalizes well to different RAG tasks.